HUBBLE SPACE TELESCOPE DAILY REPORT #1337 PERIOD COVERED: 0000Z (UTC) 02/28/95 - 0000Z (UTC) 03/01/95 Daily Status Report as of 060/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 5574 (WF/PC-2 Cycle 4: Polarization Proposal) The WF/PC-2 was used to make a visual flat calibration observation. The observation was completed as planned, and no anomalies were reported. 1.3 Completed WF/PC-2 5370 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations as part of a medium-deep sky survey. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were completed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 5234 (Faint Clusters of Galaxies (WC09): Cycle 4) The WF/PC-2 was used to make observations of faint galaxy cluster GAL-CLUS16 to monitor the observations and assess the data quality. HSTAR #5139 reported that FGS #3 failed to acquire the designated guide star in Fine Lock at 059/1410Z due to a step limit being exceeded. Coarse Track back-up was successful. No other anomalies were noted. 1.6 Completed GHRS/1 5276 (GHRS Side 1 Detector Calibration : Cycle 4) The High Resolution Spectrograph (Side One) was used to conduct a standard detector test that will be used as part of the routine calibration for the HRS. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 FGS #3 acquisition failed to Coarse Track (HSTAR 5139) at 059/1410Z. (See Note 1.5) Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Significant HSTARs: None 2.4 Operations Notes: Aside from the FGS acquisition failure to coarse track mentioned in section 2.1 and the recurrent data inversions at STGT, there were no reported operational problems. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-